A bowl of jelly beans contains 5 pink ones, 2 yellow ones, 2 green ones, 6 white ones, and 9 purple ones. If you randomly select one jelly bean, what is the probability that you get a white one?
Answer #1

Solution :

A bowl of jelly beans contains = 5 + 2 + 2 + 6 + 9 = 24

The probability that you get a white one = 6 / 24 = 0.25